# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v1

@@
Abstract
Snowflake’s Cortex AISQL is a production SQL engine that integrates native semantic operations directly into SQL. This integration allows users to write declarative queries that combine relational operations with semantic reasoning, enabling them to query both structured and unstructured data effortlessly. However, making semantic operations efficient at production scale poses fundamental challenges. Semantic operations are more expensive than traditional SQL operations, possess distinct latency and throughput characteristics, and their cost and selectivity are unknown during query compilation. Furthermore, existing query engines are not designed to optimize semantic operations.
@@ Abstract
The AISQL query execution engine addresses these challenges through three novel techniques informed by production deployment data from Snowflake customers. First, AI-aware query optimization treats AI inference cost as a first-class optimization objective, reasoning about large language model (LLM) cost directly during query planning to achieve 2–8 speedups. Second, adaptive model cascades reduce inference costs by routing most rows through a fast proxy model while escalating uncertain cases to a powerful oracle model, achieving 2–6 speedups while maintaining 90–95% of oracle model quality. Third, semantic join query rewriting lowers the quadratic time complexity of join operations to linear through reformulation as multi-label classification tasks, achieving 15–70 speedups with often improved prediction quality. AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding.
1 Introduction
@@
The growth of large language models (LLMs) [brown2020language, ouyang2022training, touvron2023llama, wei2022chain] has transformed how organizations interact with data. Modern enterprises now store vast volumes of unstructured content – documents, images, audio, and text – alongside traditional structured data in relational tables. While contemporary data warehouses excel at processing structured data with SQL, they lack native support for semantic reasoning over unstructured information. As a result, users who need to filter documents by topic relevance, join tables based on semantic similarity, or summarize customer feedback must export data to external systems, write custom scripts, or build complex pipelines that orchestrate LLM APIs outside the database. The result is significant friction: data must be moved across systems, increasing latency and cost; manual orchestration becomes error-prone; and the database loses the ability to optimize the workload end-to-end.
@@
We present Snowflake’s Cortex AISQL, a SQL engine that enables users to talk to their data, both structured and unstructured, through native AI operators integrated directly into SQL. AISQL extends SQL with primitive operators that bring LLM capabilities into the relational model: AI_COMPLETE for text generation, AI_FILTER for semantic filtering, AI_JOIN for semantic joins, AI_CLASSIFY for categorization, and AI_AGG and AI_SUMMARIZE_AGG for semantic aggregations and text summarizations, respectively. These operators naturally compose with traditional SQL constructs, allowing users to write declarative queries that blend relational operations with semantic reasoning. For example, a product manager analyzing customer feedback can filter support transcripts for conversations where customers expressed frustration (AI_FILTER), semantically join them with a product catalog to identify which specific products were discussed (AI_JOIN), classify the severity of each issue (AI_CLASSIFY), and generate executive summaries grouped by product category (AI_SUMMARIZE_AGG) – all within a single declarative SQL query that combines structured sales data with unstructured conversational text.
@@
AISQL builds on an emerging body of work in semantic query processing. Systems like LOTUS [patel2025semantic] introduced semantic operators for dataframe processing, Palimpzest [liu2025palimpzest] provided a declarative framework for LLM-powered data transformations, while ThalamusDB [jo2025thalamusdb] explored multimodal query processing with natural language predicates. Commercial systems have also adopted this direction: Google BigQuery [fernandes2015bigquery] provides AI functions like AI.CLASSIFY for semantic operations, while Azure SQL exposes row-level LLM invocations within SQL queries. However, making semantic operators efficient at production scale in a distributed database remains a fundamental challenge. Such operators are orders of magnitude more expensive than traditional SQL operations. For example, a single AI_FILTER predicate applied to a million-row table could invoke an LLM thousands of times, leading to prohibitive execution costs and latency. Traditional query optimization heuristics, such as minimizing join costs or pushing filters below joins, can produce extremely inefficient plans when AI operators are involved. Furthermore, semantic operators behave as black boxes to the query optimizer. Unlike conventional predicates, their cost and selectivity cannot be inferred from historical statistics or column distributions. Estimating how many rows will pass an AI-based filter, or how expensive that evaluation will be, is nearly impossible at compile time. Yet, these placement and cost estimates can alter overall query performance by orders of magnitude.
@@
The AISQL query execution engine addresses these challenges through a combination of novel optimization techniques and runtime adaptations, showcasing that semantic operations can be effectively integrated into database engines thus enabling use cases that analyze both structured and unstructured data. Production deployment data from Snowflake customers directly shaped our design choices. We observed that LLM inference costs far exceed traditional query processing expenses, with AI operations accounting for the dominant share of execution budgets. Multi-table queries are also prevalent, representing nearly 40% of all AISQL workloads and consuming well over half of aggregate computation time (see Section 4 for details). These observations guide the design of our optimization strategies that are applied at both the query and individual operator level:
1.
@@
AI-aware query optimization. The AISQL optimizer treats AI inference cost as a first-class optimization objective. Rather than optimizing solely for traditional metrics like join cardinality, the optimizer considers the monetary and computational cost of LLM invocations when determining operator placement and predicate evaluation order. For instance, the optimizer may pull expensive multi-modal AI_FILTER predicates above joins or reorder multiple AI predicates based on their relative cost, even when traditional selectivity-based heuristics would suggest otherwise. These optimizations enhance AISQL query performance, with the degree of improvement depending on factors like the selectivities of various query operators (e.g., predicates, joins). Across synthetic data sets and queries modeled after actual Snowflake customer workloads, we have measured 2–8 more performant execution plans. During execution, runtime statistics about predicate cost and selectivity enable dynamic reordering to adapt to actual data distributions. By reasoning about LLM cost directly during query planning, AISQL now makes even complex multi-modal workloads efficient and scalable.
2.
@@
Adaptive model cascades. To reduce inference costs and maintain query prediction accuracy, AISQL implements model cascades for AI_FILTER operations. A fast, inexpensive proxy model processes most rows, while a powerful oracle model handles only uncertain cases. The system learns confidence thresholds online through importance sampling and adaptive threshold estimation, partitioning rows into accept, reject, and uncertainty regions. Our algorithm is tailored for a distributed execution environment and achieves 2–6 speedups across various benchmarks while maintaining 90–95% of oracle model quality. On the NQ dataset, the cascade achieves 5.85 speedup and preserves the near-identical F1 score; see Section 6.2 for details.
3.
@@
Query rewriting for semantic joins. Naive semantic joins using AI_FILTER typically require a quadratic number of inference (LLM) calls making them impractical for large datasets. The AISQL engine rewrites certain join patterns into multi-label classification problems using AI_CLASSIFY, reducing complexity from quadratic to linear. Our implementation not only achieves 15–70 speedups across a number of datasets but often improves prediction quality by enabling better comparative reasoning. On the CNN dataset, the rewrite achieves 69.52 speedup, reducing execution time from 4.4 hours to 3.8 minutes; see Section 6.3 for details.
@@
The remainder of this paper is organized as follows. Section 2 describes Snowflake’s architecture with the Cortex Platform for AI workloads. Section 3 introduces AISQL’s operator semantics and SQL interface. Section 4 analyzes customer workload characteristics. Section 5 presents the query execution engine, including AI-aware query optimization, adaptive model cascades, and query rewriting for semantic joins. Section 6 provides experimental evaluation. We conclude with future work in Section 8.
2 Architecture
In this section, we discuss how Snowflake’s architecture has evolved in order to support AI workloads. Separation of compute and storage has been a key design decision in Snowflake architecture that allows users to scale the query processing resources (i.e. virtual warehouses) independently of the amount of data stored in Snowflake tables. The Cloud Services layer provides the services needed to authenticate, compile and coordinate the execution of SQL workloads.
@@ 2 Architecture
The commoditization of AI, primarily via the explosive growth and adoption of large language models (LLMs), has led to the introduction of new types of workloads as well as the evolution of existing data management workloads. AI agents today need to operate on data that are traditionally stored and managed by a warehouse. At the same time, users are interested in executing analytical workloads and collect insights from unstructured data (e.g. documents, images, audio and video); Section 4 presents an analysis of the AI workloads at Snowflake.
@@
In order to support these workloads, the Snowflake architecture evolved by adding the Cortex Platform component; the new architecture can be seen in Figure 1. The Cortex Platform is a multi-tenant component that allows the execution of both interactive (e.g. via REST) and batch AI workloads (e.g. via SQL). The Cortex Platform contains a number of components, the most important of which are: a) the Inference Engines, b) the Scheduler, and c) the API Service.
@@
The Inference Engines provide the core capabilities of the Cortex Platform. An Inference Engine is a specialized component that executes trained models (e.g. LLMs) in order to generate predictions or insights from data. An Inference Engine manages both the hardware (e.g. GPUs) and the inference software (e.g. vLLM). The Cortex Platform automatically scales the number of Inference Engines in order to support fluctuating user demand for AI inference.
@@
The Scheduler is the component responsible for orchestrating requests and assigning them to the most appropriate Inference Engine. For instance, an inference request for a specific LLM (e.g. Llama 3.1 70B) will be redirected to an Inference Engine that already hosts that model and is ready to support this request.
@@
The API Service acts as the front-end for the Cortex Platform. It receives inference requests from either the Cloud Services or the Query Processing layer, applies any API-specific business logic to the request, and then forwards the request to the Scheduler.
@@
The Cortex Platform supports a large number of models from various partners like OpenAI, Anthropic and Meta, to name a few, thus allowing users to easily build and evaluate AI applications. For each inference request, the Cortex Platform decides whether it can execute it using one of its Inference Engines or it needs to forward it to a partner endpoint. For instance, inference requests on the GPT models are forwarded to OpenAI’s endpoints by the Cortex Platform.
3 AISQL Operators
@@
In this section, we present the core operators that enable semantic data processing in Snowflake AISQL. These operators extend traditional SQL with natural language capabilities, allowing users to express complex semantic operations directly within their queries. We describe the function signatures, semantics, and usage patterns for each operator, demonstrating how they enable data-level parallelism and compose with standard SQL constructs to support rich analytical workflows.
Snowflake AISQL demonstrates a deployed system for complex semantic SQL operations using natural language in a declarative SQL interface. We support primitive operators for performing MAP, FILTER, JOIN, CLASSIFY, and REDUCE in native SQL. These operators can be composed to create rich and complex queries supported by LLMs and traditional operators for statistical analysis using JOIN or GROUP BY.
3.1 Map
@@
For example, users may query the system as follows,
@@
Each row can be processed independently, providing opportunities for data-level parallelism.
3.1.1 File Objects for Multimodal Data
To support multimodal inputs such as images, audio, or documents, AISQL provides the TO_FILE function, which references files stored in Snowflake stages. This function takes a stage path and filename, constructing a reference that can be passed to multimodal-capable models.
For example, to analyze images,
3.2 Filter
is a variation of that guarantees that the result of each row is a boolean, enabling usage in a SQL WHERE clause. For example, users can filter their queries using a natural language filter predicate as follows,
3.3 Join
Using the PROMPT object, the can be extended to multiple table and column arguments for usage in a SQL JOIN. For example, users can join multiple tables using natural language as follows,
A naive implementation applies LLM invocations joining tables with and rows, though optimization techniques help reduce this growth rate in practice.
3.4 Classify
The operator projects each row to a discrete category selected from a finite candidate set based on a natural language instruction . In contrast to , which freely generates text, constrains the output to one of the predefined categories , effectively performing a supervised classification in natural language.
Users may query the system as follows:
Example result:
| Review | AI_CLASSIFY |
| The design is beautiful and performance exceeded expectations. | positive |
| The product was fine but shipping took too long. | neutral |
| The battery stopped working after two days. | negative |
Similar to , each classification is computed independently, supporting distributed execution. The output column may then be used as an ordinary categorical attribute in downstream SQL operations such as GROUP BY, FILTER, or AGG clauses. For example, users may compute aggregate metrics per sentiment category:
@@ 3.4 Classify
3.5 Reduce
The and AI_SUMMARIZE_AGG operators reduce a column of text values to a single aggregate result. Unlike row-level operators, aggregation presents a unique challenge: tables and columns often contain more text than can fit in a single model’s context window. To address this, we employ a hierarchical aggregation strategy that processes data in batches and recursively combines intermediate results.
@@ 3.5 Reduce
Users are provided two aggregate function APIs: AI_SUMMARIZE_AGG and .
3.5.1 Summarization
Users can summarize a string expression as follows,
Example result:
| AI_SUMMARIZE_AGG |
| Opinions on the ad design are mixed, with some finding it visually appealing and others too loud or cluttered. Some suggest adding more visuals, customer testimonials, or reviews. |
The aggregate function iterates over rows, accumulating text rows in to a row buffer . Once the row buffer has exceed the BATCH_SIZE token limit, generates intermediate states containing the most important information for summarization, which are all inserted into an intermediate state buffer .
@@ 3.5.1 Summarization
3.5.2 Task Specific Aggregation
User can provide a more specific task instruction for the aggregation as follows,
Example result:
| AI_AGG |
| Most Negative Aspect: The customer’s perception that the ad is insincere and overly aggressive in its marketing approach, causing a negative physical reaction. Recommendations: 1. **Authenticity in marketing**: Ensure that marketing campaigns are genuine and align with the brand’s values and customer expectations. Avoid forced associations that may come across as insincere… |

### v2

@@
Abstract.
@@
Integrating AI’s semantic capabilities directly into SQL is a key goal, as it would enable users to write declarative queries that seamlessly blend relational operations with semantic reasoning. This would allow for effortless querying of both structured and unstructured data. However, achieving production-scale efficiency for semantic operations presents significant hurdles. Semantic operations inherently cost more than traditional SQL operations and possess different latency and throughput characteristics, making their application at large scales prohibitively expensive. Moreover, current query execution engines are not designed to optimize semantic operations, often leading to suboptimal query plans. This optimization challenge is compounded because the cost and selectivity of semantic operations are typically unknown at the time of compilation. Snowflake’s Cortex AISQL is a query execution engine that addresses these challenges through three novel techniques informed by production deployment data from Snowflake customers. First, AI-aware query optimization treats AI inference cost as a first-class optimization objective, reasoning about large language model (LLM) cost directly during query planning to achieve 2–8 speedups. Second, adaptive model cascades reduce inference costs by routing most rows through a fast proxy model while escalating uncertain cases to a powerful oracle model, achieving 2–6 speedups while maintaining 90–95% of oracle model quality. Third, semantic join query rewriting lowers the quadratic time complexity of join operations to linear through reformulation as multi-label classification tasks, achieving 15–70 speedups with often improved prediction quality. AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding.
1. Introduction
@@
The growth of large language models (LLMs) (brown2020language; ouyang2022training; touvron2023llama; wei2022chain) has transformed how organizations interact with data. Modern enterprises now store vast volumes of unstructured content – documents, images, audio, and text – alongside traditional structured data in relational tables. While contemporary data warehouses excel at processing structured data with SQL, they lack native support for semantic reasoning over unstructured information. As a result, users who need to filter documents by topic relevance, join tables based on semantic similarity, or summarize customer feedback must export data to external systems, write custom scripts, or build complex pipelines that orchestrate LLM APIs outside the database. The result is significant friction: data must be moved across systems, increasing latency and cost; manual orchestration becomes error-prone; and the database loses the ability to optimize the workload end-to-end.
@@
We present Snowflake’s Cortex AISQL, a SQL engine that enables users to talk to their data, both structured and unstructured, through native AI operators integrated directly into SQL. AISQL extends SQL with primitive operators that bring LLM capabilities to the relational model: AI_COMPLETE for text generation, AI_FILTER for semantic filtering, AI_JOIN for semantic joins, AI_CLASSIFY for categorization, and AI_AGG and AI_SUMMARIZE_AGG for semantic aggregations and text summarizations, respectively. These operators naturally compose with traditional SQL constructs, allowing users to write declarative queries that blend relational operations with semantic reasoning. For example, a product manager analyzing customer feedback can filter support transcripts for conversations where customers expressed frustration (AI_FILTER), semantically join them with a product catalog to identify which specific products were discussed (AI_JOIN), classify the severity of each issue (AI_CLASSIFY), and generate executive summaries grouped by product category (AI_SUMMARIZE_AGG) – all within a single declarative SQL query that combines structured sales data with unstructured conversational text.
@@
AISQL builds on an emerging body of work in semantic query processing. Systems such as LOTUS (patel2025semantic) introduced semantic operators for processing pandas-like dataframes, Palimpzest (liu2025palimpzest) provided a declarative framework for LLM-powered data transformations, while ThalamusDB (jo2025thalamusdb) explored multimodal query processing with natural language predicates. Commercial systems have also adopted this direction: Google BigQuery (fernandes2015bigquery) provides AI functions such as AI.CLASSIFY for semantic operations, while Azure SQL exposes row-level LLM invocations within SQL queries. However, making semantic operators efficient at production scale in a distributed database remains a fundamental challenge. Such operators are orders of magnitude more expensive than traditional SQL operations. For example, a single AI_FILTER predicate applied to a million-row table could invoke an LLM thousands of times, leading to prohibitive execution costs and latency. Traditional query optimization heuristics, such as minimizing join costs or pushing filters below joins, can produce extremely inefficient plans when AI operators are involved. Furthermore, semantic operators behave as black boxes to the query optimizer. Unlike conventional predicates, their cost and selectivity cannot be inferred from historical statistics or column distributions. Estimating how many rows will pass an AI-based filter, or how expensive that evaluation will be, is nearly impossible at compile time. Yet, these placement and cost estimates can alter overall query performance by orders of magnitude.
@@
The AISQL query execution engine addresses these challenges through a combination of novel optimization techniques and runtime adaptations, showcasing that semantic operations can be effectively integrated into database engines, thus enabling use cases that analyze both structured and unstructured data. Production deployment data from Snowflake customers directly shaped our design choices. We observed that LLM inference costs far exceed traditional query processing expenses, with AI operations accounting for the dominant share of execution budgets. Multi-table queries are also prevalent, representing nearly 40% of all AISQL workloads and consuming well over half of aggregate computation time (see Section 4 for details). These observations guide the design of our optimization strategies that are applied at both the query and individual operator level:
@@
AI-aware query optimization. The AISQL optimizer treats AI inference cost as a first-class optimization objective. Rather than optimizing solely for traditional metrics like join cardinality, the optimizer considers the monetary and computational cost of LLM invocations when determining operator placement and predicate evaluation order. For instance, the optimizer may pull expensive multi-modal AI_FILTER predicates above joins or reorder multiple AI predicates based on their relative cost, even when traditional selectivity-based heuristics would suggest otherwise. These optimizations enhance the performance of the AISQL query, with the degree of improvement depending on factors such as the selectivities of various query operators (e.g., predicates, joins). Across synthetic datasets and queries modeled after actual Snowflake customer workloads, we have measured 2–8 more performant execution plans. During execution, runtime statistics about predicate cost and selectivity enable dynamic reordering to adapt to actual data distributions. By reasoning about LLM cost directly during query planning, AISQL now makes even complex multi-modal workloads efficient and scalable.
@@
Adaptive model cascades. To reduce inference costs and maintain query prediction accuracy, AISQL implements model cascades for AI_FILTER operations. A fast and inexpensive proxy model processes most rows, while a powerful oracle model handles only uncertain cases. The system learns confidence thresholds online through importance sampling and adaptive threshold estimation, partitioning rows into accept, reject, and uncertainty regions. Our algorithm is tailored for a distributed execution environment and achieves 2–6 speedups across various benchmarks while maintaining 90–95% of oracle model quality. On the NQ dataset, the cascade achieves 5.85 speedup and preserves the near-identical F1 score; see Section 6.2 for details.
@@
Query rewriting for semantic joins. Naive semantic joins using AI_FILTER typically require a quadratic number of inference calls making them impractical for large datasets. The AISQL engine rewrites certain join patterns into multi-label classification problems using AI_CLASSIFY, reducing complexity from quadratic to linear. Our implementation not only achieves 15–70 speedups across a number of datasets but often improves prediction quality by enabling better comparative reasoning. On the CNN dataset, the rewrite achieves 69.52 speedup, reducing execution time from 4.4 hours to 3.8 minutes; see Section 6.3 for details.
@@
The remainder of this paper is organized as follows. Section 2 describes Snowflake’s architecture with the Cortex Platform for AI workloads. Section 3 introduces AISQL’s operator semantics and SQL interface. Section 4 analyzes customer workload characteristics. Section 5 presents the query execution engine, including AI-aware query optimization, adaptive model cascades, and query rewriting for semantic joins. Section 6 provides an experimental evaluation. Section 7 discusses related work. We conclude in Section 8.
2. Architecture
In this section, we describe how Snowflake’s architecture has evolved to support AI workloads. Separation of compute and storage is a key design decision in Snowflake’s architecture that lets users scale query-processing resources (i.e., virtual warehouses) independently of the amount of data stored in Snowflake tables. The Cloud Services layer handles authentication, query compilation, and coordination of SQL execution.
@@ 2. Architecture
The commoditization of AI, driven by the rapid growth and adoption of LLMs, has introduced new workloads and reshaped existing data-management tasks. AI agents today need to operate on data traditionally stored and managed in a warehouse. At the same time, users want to run analytical workloads and extract insights from unstructured data (e.g., documents, images, audio, and video). Section 4 analyzes AI workloads at Snowflake.
@@
To support these workloads, we introduced the Cortex Platform into the Snowflake architecture (Figure 1). The Cortex Platform is a multi-tenant service that executes both interactive (e.g., via REST) and batch AI workloads (e.g., via SQL). Its primary components are (a) Inference Engines, (b) the Scheduler, and (c) the API Service.
@@
The Inference Engines provide the core capabilities of the Cortex Platform. Each engine is a specialized service that runs trained models (e.g., LLMs) to generate predictions or insights from data. An engine manages both the underlying hardware (e.g., GPUs) and the inference stack (e.g., vLLM). The Cortex Platform automatically scales engines up or down to match fluctuations in inference demand.
@@
The Scheduler is the component responsible for orchestrating requests and assigning them to the most appropriate Inference Engine. For instance, an inference request for a specific LLM (e.g., Llama 3.1 70B) is routed to an engine that already hosts that model and is ready to serve it.
@@
The API Service acts as the front-end for the Cortex Platform. It receives inference requests from either the Cloud Services layer or the Query Processing layer, applies API-specific business logic, and forwards the request to the Scheduler.
@@
The Cortex Platform supports a broad set of models from partners including OpenAI, Anthropic, and Meta, enabling users to build and evaluate AI applications easily. For each inference request, the Cortex Platform determines whether to execute it using one of its Inference Engines or forward it to a partner endpoint. For instance, requests for GPT models are routed to OpenAI’s endpoints.
3. AISQL Operators
@@
In this section, we introduce the core operators that power semantic data processing in Cortex AISQL. These operators extend SQL with natural-language semantics, allowing users to express complex semantic tasks directly within queries. For each operator, we specify its signature, behavior, and typical usage, and show how the operators compose with standard SQL while exposing data-parallel execution to support rich analytical workflows. Table 1 shows the definitions of the key semantic operators implemented in AISQL.
3.1. Map
@@
For example, users may query the system as follows:
@@
3.2. Filter
is a boolean operator, enabling usage in a SQL WHERE clause. For example, users can filter the rows of a table using a natural language predicate as follows:
3.3. Join
Using the PROMPT object, the can be extended to multiple table and column arguments for use in a SQL JOIN. For example, users can join multiple tables using natural language as follows:
A naive implementation applies AI_FILTER invocations joining tables with and rows respectively, though optimization techniques help reduce this growth rate in practice (see Section 5.3).
3.4. Classify
The operator projects each row into a discrete category selected from a finite candidate set based on a natural language instruction . In contrast to , which freely generates text, constrains the output to one of the predefined categories , effectively performing a supervised classification in natural language.
Similarly to , each classification is computed independently, supporting distributed execution. The output column may also be used as an ordinary categorical attribute in downstream SQL operations such as GROUP BY, FILTER, or AGG clauses. For example, users may compute aggregate metrics per sentiment category as follows:
@@ 3.4. Classify
3.5. Reduce
AISQL supports two aggregate functions, AI_SUMMARIZE_AGG and that reduce a column of text values to a single aggregate result. For example, users can summarize the values of a string column as follows:
Unlike row-level operators, aggregation presents a unique challenge: tables and columns often contain more text than can fit in a single model’s context window. To address this, we employ a hierarchical aggregation strategy that processes data in batches and recursively combines intermediate results.
@@ 3.5. Reduce
As can be seen in the pseudo-code of Algorithm 1, the aggregate function iterates over rows, accumulating text rows in a row buffer . Once the row buffer has exceed the BATCH_SIZE token limit, generates intermediate states containing the most important information for summarization, which are all inserted into an intermediate state buffer .
@@
3.5.1. Task Specific Aggregation